OpenStudy (twizttiez):

When 40 is added to the number of miles Karen ran last week, the result is the same as adding 10 to 4 times the number of miles she ran last week. How many miles did Karen run last week?

OpenStudy (twizttiez):

@KamiBug

OpenStudy (kamibug):

This scenario can be written as: 40 + x = 10 + 4x Can you solve that? :)

OpenStudy (twizttiez):

Well 14 then idk it makes no sense to me :|

OpenStudy (kamibug):

Lol, it's okay, let's go step by step. You've gotta combine like terms, which have the same variable. So x and 4x are like terms. We wanna isolate x so let's subtract x from both sides. 40 + x - x = 10 + 4x - x Simplify? :)

OpenStudy (twizttiez):

10x??

OpenStudy (kamibug):

How'd you get it? ^_^

OpenStudy (twizttiez):

I did whatcha told me to do lol

OpenStudy (twizttiez):

You've gotta combine like terms, which have the same variable. So x and 4x are like terms. We wanna isolate x so let's subtract x from both sides. 40 + x - x = 10 + 4x - x

OpenStudy (kamibug):

Haha xD alright so x - x = 0 and 4x - x = 3x right? That gives us now: 40 = 10 + 3x Subtract 10 from both sides...

OpenStudy (twizttiez):

Still comes out to ten lol

OpenStudy (kamibug):

Lol yeah, x is finally 10. 40-10=10-10+3x 30=3x Divide both sides by 3. 10=x

OpenStudy (twizttiez):

Could you help with 3 more?

OpenStudy (kamibug):

Sure :D

OpenStudy (kamibug):

I owe you for last time Lol

OpenStudy (twizttiez):

Yay Big hug!

OpenStudy (twizttiez):

Which numbers belong to the range of the relation {(–2, 0), (1, 3), (–2, 3), (1, 0)}? Choose all answers that are correct. A. –2 B. 0 C. 1 D. 3

OpenStudy (twizttiez):

I think it's A,C

OpenStudy (kamibug):

Range is the y values and domain is the x values. A good way to remember that is that x comes before y in the alphabet and d comes before r. So here it's asking for range, which is the y values in (x,y) coordinates. Check again :)

OpenStudy (twizttiez):

So C,D

OpenStudy (kamibug):

Yup!

OpenStudy (twizttiez):

Yay nother hug! two more :)

OpenStudy (kamibug):

Lol okiedokey :3

OpenStudy (twizttiez):

Evaluate y = x – 11 when x = –1.

OpenStudy (kamibug):

What do you think you gotta do? :)

OpenStudy (twizttiez):

I dunno >_<

OpenStudy (kamibug):

Well, we're told that x = -1 right? :D So you can just plug that given value in for x! y = (-1) - 11 Got it now?

OpenStudy (twizttiez):

-12

OpenStudy (kamibug):

Perfect. ^_^

OpenStudy (twizttiez):

Yay 1 more!

OpenStudy (kamibug):

Kk

OpenStudy (twizttiez):

What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4? y = ??? x + ???

OpenStudy (twizttiez):

Huh?

OpenStudy (twizttiez):

Did you delete your response